1. You (many years ago) are deciding whether to upgrade your company's equipment from typewriters to computer technology at a cost of $2,500 per machine, and machines will be replaced one-to-one. The typewriter technology was able to output $8,000 per unit per year, and you expect the computers to be able to output $9829 per unit per year (forever), but render the typewriters completely worthless. What would the NPV be of making this upgrade, per upgraded unit? Assume a 12.6% interest rate.

2. Your production of widgets is governed by an average production cost equation, avg. cost = $1 + $5,000/(x+1000) where x is the number of goods produced. If you are currently selling 5,000 units at a price $3.08 per widget, what is the marginal cost of selling an additional unit?
Answer #1

1.

=-2500+(9829-8000)/12.6%=12015.87

2.

Total cost=x+5000*x/(x+1000)

Marginal cost=1+5000*(0+1000*(x+1000)^(-2))=1+5000*1000/(x+1000)^2=5.97